# FaceEditTalker: Interactive Talking Head Generation with Facial Attribute Editing

### v1

@@ 3.2 Task Formulation
3.3 Image Feature Space Editing Module
The semantic encoder $E_{\text{sem}}$ extracts global facial semantics from the input image, encoding them into low-dimensional vectors akin to StyleGAN’s style vectors, enabling linear transformations for attribute editing. Given a semantic code $z_{\text{sem}}$ and an attribute direction vector $w_{\text{attr}}$, attribute manipulation is expressed as:
$z^{\prime}_{\text{sem}}=z_{\text{sem}}+\alpha\cdot w_{\text{attr}},$ (1)
where $\alpha$ controls the intensity of the change. To enable text-driven editing and align more clearly with the illustrated process, text input is processed in conjunction with an attribute library by a linear classifier to compute the attribute direction vector $w_{\text{attr}}$. This $w_{\text{attr}}$ is then added to the original semantic code $z_{\text{sem}}$ according to Eq (1) to obtain the edited semantic code $z^{\prime}_{\text{sem}}$. A reconstruction loss $L_{\text{rec}}$ ensures that non-target regions remain consistent:
$L_{\text{rec}}=|z^{\prime}-z_{\text{static}}|.$ (2)
The stochastic encoder, designed as a Unet-based diffusion model [47]. First use the $E_{img}$ encode the input image into an initial latent representation $Z_{0}$. This representation undergoes a noise diffusion process to capture fine details, resulting in the final latent representation $Z_{T}$. We refer to this complete encoding and diffusion procedure as the stochastic encoder. The forward process, conditioned on $z_{\text{sem}}$, is defined as:
$z_{t+1}=\sqrt{\alpha_{t+1}}f_{\theta}(z_{t},t,z_{\text{sem}})+\sqrt{1-\alpha_{t+1}}\epsilon_{\theta}(z_{t},t,z_{\text{sem}}).$ (3)
This reverse diffusion effectively complements the semantic code’s lack of detail. During video generation, the semantic code $z_{\text{sem}}$ manages global attribute control, while the stochastic code $Z_{T}$ preserves local detail consistency, ensuring smooth and realistic facial transformations across frames.
3.4 Audio-Driven Video Generation Module
This module generates video using audio input, semantic code, and stochastic code. The audio input is first processed to extract audio features $A^{1:T}$. These features are then transformed into an audio-driven facial landmark sequence $L^{1:T}$ through a landmark prediction network, represented as:
$L^{1:T}:(l^{1},l^{2},\dots,l^{T})=P_{ldm}(A^{1:T}:(a^{1},a^{2},\dots,a^{T})),$ (4)
where $P_{ldm}$ encapsulates the transformation from audio features to landmarks, involving a pre-trained wav2vec model and a regression model.
Subsequently, the facial landmark sequence $L^{1:T}$ is processed by a landmark feature extractor to obtain the corresponding landmark features $F^{1:T}$:
$F^{1:T}:(f^{1},f^{2},\dots,f^{T})=E_{ldm}(L^{1:T}:(l^{1},l^{2},\dots,l^{T})),$ (5)
where $E_{ldm}$ represents the landmark feature extractor. It adopts multi-scale strategies and cross-attention mechanisms to capture multi-level dynamics and enhance feature correlations, improving precision and temporal consistency in video generation.
The extracted facial landmark features $F^{1:T}$ is fused with the stochastic code $Z_{T}$ to produce the feature $K$. During the sampling process of the diffusion model, we employ an innovative conditional feature injection strategy, where the high-level semantic information $z_{\text{sem}}$ is injected as a global facial attribute control signal into the conditional diffusion model, while the target feature $K$ serves as dynamic motion information. This mechanism ensures stable expression of semantic attributes alongside synchronized audio-driven facial movements:
$z_{t-1}=\sqrt{\alpha_{t}}z_{t}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}(z_{t},z_{\text{sem}},K,t),$ (6)
where $\alpha_{t}$ represents the noise scheduling parameters, and $\epsilon_{\theta}$ is the conditional denoising network that guides the denoising process based on the latent variable $z_{\text{sem}}$ and the target features $K$.
Finally, the denoised latent variable $z_{0}$ is decoded into a sequence of video frames, which not only exhibit dynamic facial expressions synchronized with the audio, but also, depending on the switch selection, achieve facial attribute editing of the speaker based on the different semantic encoding $z_{\text{sem}}$, resulting in a high-quality editable talking head video.
3.5 Training and Inference Pipeline
Training process: This model is divided into three stages. First stage jointly trains the semantic encoder and random encoder to extract semantic and random detail features of images by minimizing the mean squared error between predicted and actual noise, with the loss function:
$\mathcal{L}_{\text{simple}}=\sum_{t=1}^{T}\mathbb{E}_{x_{o},\epsilon_{t}}[||\epsilon_{\theta}(x_{t},t,z_{\text{sem}})-\epsilon_{t}||_{2}^{2}].$ (7)
@@ 3.5 Training and Inference Pipeline
Second stage trains an image semantic linear classifier, optimizing the model using cross-entropy loss to accurately classify image attributes, with the loss function:
$\mathcal{L}=-\frac{1}{N}\sum_{l=1}^{N}[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})].$ (8)
Third stage trains the latent space diffusion model generation network, fusing semantic and key point features, and generating images by minimizing the mean squared error between predicted and actual noise, with the loss function:
$\mathcal{L}=\sum_{t=1}^{T}\mathbb{E}_{z_{o},\epsilon_{t}}[||\epsilon_{\theta}(z_{t},t,z_{\text{sem}},z_{l})-\epsilon_{t}||_{2}^{2}].$ (9)
Inference process: First, face key points and pose information are extracted from the audio sequence and reference image, then combines semantic encoding and attribute editing information and inputs them into the trained latent space diffusion model to finally generate the video sequence of a talking head.
4 Experiments
4.1 Experimental Settings
@@
Datasets. We trained the encoder of the dual-layer latent architecture using the FFHQ dataset [22], which offers high-resolution facial images with diverse attributes including age, race, expression, facial structure, hair features, and accessories, ideal for learning complex feature representations. For the linear classifier, we used the CelebA-HQ dataset [26] with binary labels for 40 facial attributes to enhance attribute feature separation and model generalization. In the audio-driven facial animation generation stage, we utilized the HDTF dataset [60], containing lip-sync videos from over 300 speakers, along with VoxCeleb2 [33] and VFHQ [52] datasets to improve the model’s ability to learn complex mappings between speech and facial movements under various environmental conditions. Additionally, we applied LatentSync [27] to refine dataset quality by resampling videos, removing those with low synchronization confidence, correcting audiovisual offsets, and filtering out clips with poor HyperIQA scores, thereby enhancing lip-sync accuracy and visual quality.
Comparison Methods. To the best of our knowledge, there is no existing method capable of generating high-resolution, audio-driven speaker videos with editable facial attributes. For a comprehensive evaluation of our proposed method, we first generate results using semantic features extracted by the high-level semantic encoding module, ensuring identity consistency with reference images. Our method is compared against several state-of-the-art (SOTA) lip synchronization approaches categorized into three groups. Wav2Lip [38] optimizes direct mappings between audio and lip motion for highly synchronized lip movements while preserving facial textures. SadTalker [58] employs explicit facial landmarks and adversarial networks to produce smooth animations. DiffTalk [46], EchoMimic [9], and Hallo [54] leverage diffusion models to model conditional distributions between audio and facial movements, achieving higher-quality talking videos and strong generalization capabilities for out-of-distribution subjects. This comparison aims to evaluate our method’s performance relative to current leading techniques in audio-driven talking head generation.
Evaluation Metrics. For evaluating our method, we employ several metrics. Image generation quality is assessed using FID [12], SSIM [3], PSNR [19], and CPBD [34]. Lip motion accuracy is evaluated with M-LMD and F-LMD [8], while Syncconf [39] measures lip movement-audio synchronization. Additionally, we edit semantic features from the dual-layer semantic encoding module using a linear classifier to produce edited video results. These are compared against state-of-the-art video editing methods such as Latent-Transformer [56], STIT [49], and Diffusion-Video-Autoencoders [24], using TL-ID and TG-ID [49] as evaluation metrics.
4.2 Comparison with Other Methods
Table 1: Quantitative evaluation of our approach compared with SOTAs. Our method achieves the best performance on FID, SSIM (second best in EchoMimic), lip sync metrics, and keypoint error metrics, demonstrating superior overall quality and synchronization accuracy.
Video Quality Lip Sync Keypoint Error
@@ 4.2 Comparison with Other Methods
Quantitative Evaluation. We quantitatively compared FaceEditTalker with existing state-of-the-art audio-to-face generation methods on the HDTF and VoxCeleb2 datasets. As shown in the table, FaceEditTalker demonstrates excellent performance in metrics like image feature similarity, structural similarity, and image quality, leveraging a latent space diffusion model. Our method achieves significant improvements compared to previous diffusion-based approaches, attributed to a more advanced model framework. Furthermore, our model shows strong SyncNet scores, benefiting from optimizations such as dataset correction, multi-scale strategies, and cross-attention mechanisms.
Table 2: Quantitative results of our approach compared with SOTAs. Our method achieves the best performance on identity consistency in face attribute editing.
@@
Furthermore, regarding identity consistency in face attribute editable talking head generation, we conducted quantitative evaluation. Our method edits semantic features to generate videos with 20 attributes, compared against video editing algorithms. Evaluation of identity consistency between frames showed that while our generative model achieved similar overall identity consistency as video editing methods, it significantly excelled in TL-ID.
Qualitative Evaluation. Figure 3 compares the generation quality of FaceEditTalker with existing advanced methods. While previous approaches often prioritize specific aspects such as lip synchronization or introduce artifacts and distortions when aiming for expressiveness, our proposed algorithm generates videos with better overall image quality and more accurately captures fine facial expression details, closely matching the original video. FaceEditTalker particularly excels at handling nuanced facial actions like eye closure and mouth opening, contributing to a higher level of realism in the generated results.
@@
Figure 4: Video generation results with the editing feature enabled. Using three different reference images and the same audio clip, we demonstrate the editing and speaker generation effects under four different attribute editing categories with various sub-attributes. More attribute editing results can be found in Appendix C and Supplementary materials.
@@
Table 3: User Study Results for Editable Facial Attribute Talking Head Generation
@@
Method Lip Sync $\uparrow$ Realism $\uparrow$ Video Quality $\uparrow$ Attribute Editing Effect $\uparrow$
Original Video 4.80 4.90 4.80 Not Supported
Wav2Lip[38] 3.60 3.10 3.70 Not Supported
SadTalker[58] 2.70 2.20 2.80 Not Supported
DiffTalk[46] 3.40 3.00 2.90 Not Supported
EchoMimic[9] 3.40 3.70 3.90 Not Supported
Hallo[54] 3.80 3.30 3.60 Not Supported
Latent-Transformer[56] 3.30 3.80 3.60 3.40
STIT[49] 3.30 3.30 3.00 3.30
Diffusion Video Autoencoders[24] 3.10 3.50 3.10 3.30
Our Method 3.80 3.60 3.70 4.40
User Study. We conducted a user study (10 people, score 1-5) on lip sync accuracy, realism, video quality and attribute editing effects using a five-point scale. Our method achieved high scores across all metrics: 3.8 for lip-sync (benefiting from multi-scale keypoint features and SyncNet preprocessing), strong performance in realism and video quality (attributed to the latent space diffusion model’s capabilities), and superior performance in attribute editing. As shown in Table 3, methods marked with ’N/A’ in the Attribute Editing Effect column indicate that they do not support this functionality.
4.3 Analysis and Ablation Study

### v2

@@ III-E Training and Inference Pipeline
where $\epsilon_{\theta}(\cdot)$ denotes the conditional denoising network that predicts the noise component, and $\epsilon_{t}$ represents the ground-truth noise at timestep $t$.
@@
The second stage involves training a linear classifier using cross-entropy loss to accurately predict image attributes:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(\hat{y}_{i,c}),$ (12)
where $N$ denotes the total number of samples, $C$ is the number of attribute categories, and $y_{i,c}$ and $\hat{y}_{i,c}$ correspond to the ground truth labels (encoded as one-hot vectors) and the predicted attribute probabilities, respectively.
The third stage involves training a conditional diffusion model in the latent space. At each diffusion step, the model receives a noisy latent vector $Z_{T}$ along with the semantic code $z_{\text{sem}}$ and motion lanmark features $F$. These conditioning signals are fused to guide the denoising network, allowing precise control over both static attributes and dynamic motion. This design facilitates stable training and enables high-quality, editable talking head synthesis. The training objective is to minimize the mean squared error between the predicted and ground-truth noise:
[…]
where the semantic and motion features are fused to condition the latent diffusion process, enabling the model to jointly perform facial attribute editing and dynamic motion synthesis within a stable and efficient latent space.
Inference process: We generate talking head videos using only an input audio clip and a reference image, with optional attribute specifications for editing. During inference, the audio and reference image are used to extract the landmark feature $F$, while the reference image also provides a semantic code $z_{sem}$. If editing is required, the semantic code is fused with the corresponding attribute vector $w_{att}$. The landmark feature $F$ is then combined with the stochastic code $Z_{T}$ and fed into the DDIM sampler, which performs denoising under the guidance of the semantic code $z_{sem}$ on the latent embedding that encodes landmark information. Finally, this process yields a temporally consistent sequence of video frames, where facial movements are synchronized with the audio and facial attributes are controllably edited as specified, resulting in high-quality, realistic, and editable talking head videos.
IV Pseudo-code for this method
To help researchers better understand the overall workflow of our approach, we present the pseudo-code along with a description of the training and inference processes.
IV-A Training and Inference Process Description
The method is organized into three principal training modules. The first two stages focus on training the Image Feature Space Module, while the third stage trains the Audio-Driven Video Generation Module, followed by the inference process.
  1. 1.
    Joint Training of Semantic Encoder and Stochastic Encoder: Face images are encoded to extract high-level semantic and stochastic features, optimized using a diffusion model.
  2. 2.
    Training the Image Semantic Linear Classifier: Attribute variation directions are learned in the high-level semantic space, enabling precise attribute classification.
  3. 3.
    Training the Audio-Driven Video Generation Module: The injected noise, together with landmark features, is incorporated into DDIM, where the generation process is guided by semantic code to produce high-quality, identity-consistent talking face videos.
  4. 4.
    Inference Process: In the inference phase, the model uses the input audio sequence, reference image, and attribute information to generate facial motion features, which are processed by the diffusion model to produce high-quality editable talking face videos.
IV-B Training Stage
Algorithm 1 Joint Training of Semantic Encoder and Stochastic Encoder
2:Output: Weight vectors corresponding to each attribute list
3:for each epoch do
4:  for each $x_{i}\in\text{FFHQ(image)}$ do
5:   $z_{\text{sem}}=\text{SemanticEncode}(x_{i})$ $\triangleright$ Semantic Encoder Forward Pass
6:   $t=\text{RandomTimeStep}()$
7:   $Z_{0}=\text{T=ImgEncoder}(x_{i})$
8:   $Z_{T}=\text{AddNoise}(Z_{0})$
9:   $\hat{x}=\hat{q}(Z_{T},z_{\text{sem}})$ $\triangleright$ Model Forward Pass
10:   $e=\text{ComputeTarget}(Z_{t})$ $\triangleright$ Compute Target
[…]
12:   Backpropagate and update parameters
13:  end for
14:  Save model at the end of each epoch
15:end for
Algorithm 2 Training the Image Attributes Linear Classifier
1:Input: Face image set with facial attributes, attribute labels, learning rate
2:Output: Weight vectors corresponding to each attribute list
3:for each epoch do
4:  for each attribute $x_{i}\in\text{Att}(x_{1},x_{2},\dots,x_{n})$ do
5:   $z_{\text{att}}=\text{SemanticEncoder}(x_{i})$
6:   $y_{\text{labels}}=\text{GetAttributeLabels}(x_{i},\text{FFHQ}(image))$
7:   $W_{att}=\text{InitializeWeightVector}()$
8:   $b=\text{InitializeBias}()$
9:   for each $x_{i}\in\text{FFHQ(image)}$ do
10:     for each $z_{\text{att}}$ do
11:      $\hat{y_{i}}=\text{Sigmoid}(W_{att}z_{\text{att}}+b)$ $\triangleright$ Forward Pass
12:      $L=\text{CrossEntropyLoss}(\hat{y_{i}},y_{i})$ $\triangleright$ Cross-Entropy Loss
13:      Backpropagate and update parameters
14:     end for
15:   end for
16:  end for
17:  Save weight vectors and attribute-label pairs $(y_{i},w_{att,i})$
18:end for
Algorithm 3 Training the Audio-Driven Video Generation Module
1:Input: reference image set, audio sequence $S_{audio}$, learning rate
2:Output: Diffusion model parameters
3:Data Preprocessing: Use Wav2Vec to extract audio feature $A^{1:T}$ sequence from audio sequence $\mathcal{S}_{audio}$
4:for each epoch do
5:  for each batch $A^{1:T},~x_{\text{ref}}$ do
6:   $Z_{0}=\text{ImgEncoder}(x_{ref})$
7:   $Z_{T}=\text{AddNoise}(Z_{0})$
8:   $L^{1:T}=LandmarkPredictor(A^{1:T})$
9:   $F^{1:T}=LandmarkFeatureExtractor(L^{1:T})$
10:   $K=Concat(F^{1:T},Z_{T})$
11:   $z_{\text{sem}}=\text{SemanticEncoder}(x_{\text{ref}})$
12:   $t=\text{RandomTimeStep}()$
13:   $\hat{\epsilon}=\epsilon_{\theta}(K,t\mid z_{\text{sem}})$
14:   $L=\text{MSE}(\hat{\epsilon},\epsilon)\triangleright$ MSE Loss
15:   Backpropagate and update parameters
16:  end for
17:  Save model at the end of each epoch
18:end for
IV-C Inference Stage
Algorithm 4 Controllable Talking Head Generation with Facial Attribute Editing
1:Input: Reference image $x_{ref}$, audio sequence $\mathcal{S}_{audio}$, attribute label-weight pairs $(y_{i},w_{\text{att,i}})$, attribute editing magnitude $\alpha$, diffusion time steps
2:Output: Speaker video frame sequence $S_{video}$
[…]
8:if $\alpha$ is not None then
9:  $z_{sem}=z_{sem}+\alpha\cdot w_{att}$
10:end if
[…]
V Experiments
V-A Experimental Settings
Data Preprocessing. During training and validation, videos were sampled at 25 frames per second (FPS) and audio at 16 kHz. To ensure data consistency, videos were cropped and resized to a resolution of $512\times 512$ pixels. Audio–video synchronization was achieved using Mel-spectrogram representations with a window length and hop length of 640 samples. In stage one, we used 16 input frames with a stride of 4; in stage three, 16 input frames with a stride of 1 and stride augmentation were employed.
Training Configuration. Our method was trained on two NVIDIA A100 GPUs. The first and second stages were trained for 100 hours and the third stage was trained for 160 hours. The main parameters are shown in Table I:
TABLE I: Experimental Parameter Settings.
Parameters Value/Range
Random Seed 0
Image Size 512*512
Batch Size 16
Learning Rate 0.0001
Training Epochs 20000
Embedding Layer Channels 512
Diffusion Timesteps 1000
@@ V-A Experimental Settings
Datasets. We trained the dual-layer latent architecture encoder using the FFHQ dataset [19], which offers high-resolution facial images with various attributes including age, race, expression, facial structure, hair features and accessories, ideal for learning complex feature representations. For the linear classifier, we used the CelebA-HQ dataset [65] with binary labels for 40 facial attributes to enhance attribute feature separation and model generalization. In the audio-driven facial animation generation stage, we utilized the HDTF dataset [66], containing lip-sync videos from more than 300 speakers, along with VoxCeleb2 [67] and VFHQ [68] datasets to improve the model’s ability to learn complex mappings between speech and facial movements under various environmental conditions. Additionally, we applied LatentSync [69] to refine dataset quality by resampling videos, removing those with low synchronization confidence, correcting audiovisual offsets, and filtering out clips with poor HyperIQA scores, thereby enhancing lip-sync accuracy and visual quality.
TABLE II: Quantitative evaluation of our approach compared with representative approaches.
Video Quality Lip-sync Keypoint Error
@@
Comparison Methods. To the best of our knowledge, there is no existing method capable of generating high-resolution, audio-driven speaker videos with editable facial attributes. For a comprehensive evaluation of our proposed method, we first generate results using semantic features extracted by the high-level semantic encoding module, ensuring identity consistency with reference images. We compare our method with several representative and widely used lip synchronization methods. Wav2Lip [1] optimizes direct mappings between audio and lip motion for highly synchronized lip movements while preserving facial textures. SadTalker [41] employs explicit facial landmarks and adversarial networks to produce smooth animations. DiffTalk [43], EchoMimic [42], and Hallo [45] leverage diffusion models to model conditional distributions between audio and facial movements, achieving higher-quality talking videos and strong generalization capabilities for out-of-distribution subjects. This comparison aims to evaluate our method’s performance relative to current leading techniques in audio-driven talking head generation.
Evaluation Metrics. For evaluating our method, we employ several metrics. Image generation quality is assessed using FID [70], SSIM [71], PSNR [72], and CPBD [73]. Lip motion accuracy is evaluated with M-LMD and F-LMD [74], while Syncconf [75] measures lip movement-audio synchronization. Additionally, we edit semantic features from the dual-layer semantic encoding module using a linear classifier to produce edited video results. These are compared against representative video editing methods such as Latent-Transformer [23], STIT [24], and Diffusion-Video-Autoencoders [26], using TL-ID and TG-ID [24] as evaluation metrics.
V-B Performance Comparison
Quantitative Evaluation. We quantitatively compared FaceEditTalker with representative audio-to-face generation methods on the HDTF and VoxCeleb2 datasets. As shown in Table II, FaceEditTalker achieves the best FID and M-LMD scores on both datasets, while maintaining competitive SSIM (second only to EchoMimic), PSNR, and lip-sync performance. This demonstrates strong overall performance across image feature similarity, structural similarity, and visual quality, enabled by the latent space diffusion model for fine-grained facial attribute control. Compared to prior diffusion-based approaches, our method benefits from a more advanced framework with enhanced semantic editing capabilities. Superior visual quality and robust audio-visual alignment, reflected in high SyncNet scores, are further supported by dataset corrections to reduce alignment errors, multi-scale modeling of global and local facial motions, and cross-attention mechanisms reinforcing audio–facial correspondence.
Despite these strengths, our method achieves moderate performance on metrics evaluating local lip articulation and fine-grained audio-visual alignment. Specifically, Min Dist indicates limited precision in modeling subtle lip movements, AVConf reflects suboptimal audio-lip synchronization, and F-LMD suggests reduced fidelity in reconstructing local lip contours. These limitations likely arise from the landmark-guided diffusion framework emphasizing global facial consistency, which can smooth localized lip variations, and the lack of fine-grained supervision in mapping audio features to lip shapes. Future work incorporating localized lip refinement, enhanced audio-lip feature mapping, and fine-grained supervisory signals could further improve lip motion fidelity and overall audio-visual coherence.
Identity Consistency Evaluation. We quantitatively evaluated identity consistency in face attribute editable talking head generation, as presented in Table III. Our method edits semantic features to generate videos with 20 attributes, compared to video editing algorithms. Evaluation of identity consistency between frames showed that while our generative model achieved similar overall identity consistency as video editing methods, it significantly excelled in TL-ID.
TABLE III: Quantitative results of our approach compared with representative approaches.
@@ V-B Performance Comparison
Figure 5: Video generation results with the editing feature enabled. Using three different reference images and the same audio clip, we demonstrate the editing and speaker generation effects under four different attribute editing categories with various sub-attributes.
Qualitative Evaluation. Fig. 4 compares the generation quality of FaceEditTalker with existing advanced methods. While previous approaches often prioritize specific aspects such as lip synchronization or introduce artifacts and distortions when aiming for greater expressiveness, our proposed algorithm generates videos with better overall image quality and more accurately captures fine facial expression details, closely matching the original video. FaceEditTalker particularly excels at handling nuanced facial actions like eye closure and mouth opening, thereby contributing to a higher level of realism in the generated results.
User Study. To comprehensively evaluate the quality of the generated talking head videos, we conducted a user study involving 50 participants, focusing on four key dimensions: lip-sync accuracy, realism, video quality, and attribute editing effects. Each dimension was rated independently using a five point Likert scale (1 = poor, 5 = excellent). To ensure fairness and consistency, all methods—including our proposed approach and several baseline systems—were applied to the same source video under identical conditions. For methods supporting facial attribute editing, we introduced a predefined modification—the addition of eyeglasses—chosen for its visual saliency and broad applicability across different models. In contrast, methods lacking editing capability produced unaltered outputs. Participants were shown the generated videos and asked to assess each quality dimension separately.
@@
TABLE IV: User Study Results for Editable Facial Attribute Talking Head Generation.
@@
Method Lip-sync $\uparrow$ Realism $\uparrow$ Video Quality $\uparrow$ Attribute Editing Effect $\uparrow$
Original Video 4.80 4.90 4.80
Lip[1] 3.82 3.38 3.46
SadTalker[41] 3.32 3.24 3.28
DiffTalk[43] 2.96 3.30 2.96
EchoMimic[42] 3.48 3.34 3.06
Hallo[45] 3.44 3.30 3.32
Latent-Transformer[23] 3.34 3.32 3.30 3.48
STIT[24] 3.08 3.18 2.96 3.04
Diffusion Video Autoencoders[26] 2.96 3.22 3.44 3.26
Our Method 3.58 3.18 3.50 4.04
In the study, lip-sync accuracy was evaluated based on lip movement alignment with speech and consistency with the original video. Our method utilizes the Wav2Vec framework to extract audio features for predicting facial mesh and keypoints. Leveraging multi-scale keypoints and SyncNet preprocessing, it achieves strong lip-sync with a mean score of 3.58, second only to Wav2Lip. Realism and video quality were judged by the naturalness of expressions and visual clarity. The realism score averaged 3.18, limited by reliance on audio features alone, which sometimes caused rigid or less expressive facial movements. Video quality scored slightly higher at 3.50, benefiting from our latent diffusion architecture. For attribute editing, participants rated perceptual clarity and controllability; our method scored 4.04, outperforming most baselines—some of which lacked editing support (‘✗’ in Table IV). Overall, as shown in Table IV, our approach surpasses most baselines, demonstrating its effectiveness in generating realistic, high-quality, and controllably editable talking head videos.
V-C Analysis and Ablation Study
Effectiveness of the Facial Landmark Feature Extractor. We conducted ablation experiments on two datasets to validate the effectiveness of the proposed components in the facial landmark feature extractor. As illustrated in Table V, I denotes the Original Video serving as the ground-truth reference, II corresponds to the Multi-layer Convolution baseline, III represents the Multi-scale Strategy, IV indicates the Multi-scale combined with Cross-Attention, and V denotes the Edited Semantic Encoding. The results demonstrate that the multi-scale strategy (III) facilitates more accurate modeling of fine-grained facial dynamics such as lip and eye motions, while the integration of cross-attention (IV) further improves all evaluation metrics, particularly AVOffset and M-LMD, thereby enhancing temporal consistency with the original video. Moreover, the Edited Semantic Encoding (V) shows minimal impact on the quantitative evaluation metrics.